OpenStudy (anonymous):

find the distance using the distance formula for the following two points A (2,8) and B (-6,-4) A 14 B 15 C 14.42 D 15.42

OpenStudy (anonymous):

\[\sqrt{(-6-2)^{2} + (-4-8)^{2}} = \sqrt{208}= 14.42\]

OpenStudy (bahrom7893):

Distance formula is: sqrt( (x2 - x1)^2 + (y2 - y1)^2 ) = sqrt( (-6 - 2)^2 + (-4 - 8)^2 ) = sqrt( (-8)^2 + (-12)^2) = sqrt ( 64 + 144 ) = sqrt (208) = 14.42

OpenStudy (anonymous):

Distance formula\[ \sqrt{(x_{1}-x_{2})^{2} +({y_1}-y_{2})^{2}}\] Insert values\[ \sqrt{(2--6)^{2} +(8--4)^{2}}\] \[= \sqrt{8^{2} +12^{2}}\] \[= \sqrt{64 +144}\] \[= \sqrt{208}= 4\sqrt{13}=14.42\] The answer is D